TEX User Groups Around the World

CSTUG (Czech and Slovak Republics)

Dr. Karel Horák, President Československé sdružení uživatelů TEXu CSTUG, c/o MÚK UK Sokolovská 83 CS-186 00 Praha 8, Czech Republic

horakk@earn.cvut.cz
ftp: ftp.muni.cz/pub/tex
Web: http://www.muni.cz/TeXhelp/TeX-homepage.html

CyrTUG (Russia)

Irina Makhovaya, Executive Director Associaciia Pol’zovateleĭ Kirillicheskogo TEX’a Mir Publishers 2, Pervyĭ Rizhskiĭ Pereulok Moscow 129820, Russia Tel: 095 286-0622, 286-1777 FAX: 095 288-9522

cyrtug@mir.msk.su

DANTE e.V. (German-speaking)

Joachim Lammarsch, President Deutschsprachige Anwendervereinigung TEX e.V. Postfach 101840 D-69008 Heidelberg, Germany Tel: +6221 29766 FAX: +6221 167906

dante@dante.de
ftp: ftp://ftp.dante.de
Web: http://www.dante.de

Estonian User Group Enn Saar, Tartu Astrophysical Observatory, Toravere EE 2444 Estonia

saar@aai.ee

GUST (Poland)

Tomasz Przechlewski (President) Polska Grupa Użytkowników Systemu TEX Centrum Fizyki Teoretycznej PAN Aleja Lotników 32/46 PL-02-668 Warszawa, Poland

e-mail:  GUST@ifpan.edu.pl
Web: http://www.cc.uni.torun.pl/GUST/GUSThp.html

GUTenberg (French-speaking)

Michel Goossens, President Groupe francophone des Utilisateurs de TEX Association GUTenberg BP 10 F-93220 Gagny principal, France Tel: +33 1 44 32 37 96 FAX: +33 1 44 32 20 80

gut@irisa.fr
Web: http://www.ens.fr/gut

ITALIC (Irish)

Not a user group yet. Public discussion list: ITALIC-L on

listserver listserv@irlearn.ucd.ie.

iTeXnici (Italian)

An unofficial group yet. Contact: Giovanni MenoZero Pensa

pensa@dsi.unimi.it.
Web: http://www.dsi.unimi.it/Users/Students/pensa/
tex/Tipo/Bibliografia.html

Lithuanian TEX Users Group Vytas Statulevicius, Chair Akademijos 4 2600 Vilnius, Lithuania Tel: +370 2 359 609 FAX: +370 2 359 804

statulevicius@mii.lt

Nordic TEX Users Group (Scandinavian countries) Dag Langmyhr, Chair Nordic TEX Users Group Department of Informatics P.O. Box 1080 Blindern University of Oslo N-0316 Oslo, Norway Tel: +47 22 85 24 50 FAX: +47 22 85 24 01

dag@ifi.uio.no
Web: http://www.ifi.uio.no/~dag/ntug/ntug.html

NTG (Dutch-speaking)

Erik Frambach, Chair Nederlandstalige TEX Gebruikersgroep Postbus 394 NL-1740 AJ Schagen, The Netherlands,

ntg@nic.surfnet.nl
Web (experimental): http://ei0.ei.ele.tue.nl/ntg

Spanish TUG (no formal group)

Not a user group yet. Electronic mail discussion list: spanish-tex@eunet.es. Send subscription requests to

listserv@eunet.es.

Julio Sanchez GMV SA Isaac Newton 11 PTM Tres Cantos E-28760 Madrid, Spain

jsanchez@gmv.es

TeXCeH (Slovenian User Group)

(part of larger group, Informatica) Vladimir Batagelj Jadranska 19 61111 Ljubljana, Slovenia

texceh@uni-lj.si

TUG (International user group)

Michel Goossens, President Patricia Monohon, Executive Director TEX Users Group 1850 Union St., Suite 1637 San Francisco, CA 94123 USA Tel: +1 415-982-8449 FAX: +1 415-982-8559

tug@tug.org

UK TUG (United Kingdom)

R. Fairbairns, Chairman UK TEX Users’ Group For information: Peter Abbott 1 Eymore Close, Selly Oak Birmingham B29 4LB, England Tel: +1 121 476 2159

uktug-enquiries@tex.ac.uk
Web: http://www.tex.ac.uk/UKTUG